OpenStudy (anonymous):

Find the number of sides in the polygon with whose angles measures have the given sum. 18,000 degrees a. 102 b. 98 c. 200 d. 120

OpenStudy (anonymous):

What is the formula for finding the number of degrees in a polygon? Do you know?

OpenStudy (anonymous):

No I don't

OpenStudy (anonymous):

Ok so the formula is \[(n-2)180=degrees of a polygon\]

OpenStudy (anonymous):

Where n= the number of sides a polygon has. Does that help?

OpenStudy (anonymous):

So 18000/180= 100-2 or +2?

OpenStudy (anonymous):

You are on the right track but that is not quite right. You have a polygon with 18,000 degrees. (n-2)180=18000 Instead of dividing by 180 we need to distribute first. 180n-360=18000 Add 360 on both sides 180n=18360 Divide 180 on both sides n=102 so n is the variable for how many sides so you have 102 sides in this polygon.

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

Holy pellet, Thank you O: